OpenStudy (anonymous):

The pair of points is on the graph on an inverse variation. Find the missing vale . (9,5) and (x,6)

OpenStudy (anonymous):

9 = k5 k = 9/5 6 = 9/5 x 10/3 =x well it is a guess that all you should confirm it

OpenStudy (anonymous):

Is the answer 9 ?

OpenStudy (anonymous):

i got 10/3

OpenStudy (a_clan):

Inverse variation, 5=k/9 k=45 6=k/x or x=k/6=45/6=15/2=7.5